18 2/3 x 4 1/6 x 3/5 = 46 2/3 Please show the steps to get to this answer.
Found 2 solutions by checkley77, edjones:
Answer by checkley77(12844)

18 2/3 x 4 1/6 x 3/5 = 46 2/3
[(18*3+2)/3][(4*6+1)/6][(3/5}=[(46*3+2)]/3
[(54+2)/3][(25+4)/6][3/5]=[(138+2)/3]
[56/3][25/6][3/5]=140/3
[(56*25*3)/90]=140/3
4200/90=140/3
140/3=140/3

Answer by edjones(8007)

18 2/3 x 4 1/6 x 3/5
=56/3 * 25/6 * 3/5
=4200/90
=140/3
=46 2/3
.
Ed